# SplitVAEs: Decentralized scenario generation from siloed data for stochastic optimization problems

### v1

@@ III-D Loss Functions
Reconstruction Loss: We employ the binary cross-entropy (BC) loss function at the $n$-th edge location for all $n\in\mathcal{N}$ to measure the mean reconstruction error between predicted values $(\tilde{y})^{n}$ and observed values $(y)^{n}$ as represented in Equation (7).
$L^{BC}_{n}=-\frac{1}{B}\sum\limits_{b=1}^{B}\Big{[}\tilde{y}^{b}log(y^{b})+(1-\tilde{y}^{b})log(1-y^{b})\Big{]}$ (7)
In Equation (7), $B,\tilde{y}_{b},y_{b}$ represent the batch size, predicted, and target data for scenario generation, respectively. To preserve the simplicity of notations, we assume without loss of generality that $L^{BC}_{n}$ is computed concerning flattened representations of local tensor data and estimates $(y)^{n},(\tilde{y})^{n}$ of dimension $d^{n}$. We note that the computation of $L^{BC}_{n}$ can be carried out independently at each edge location since it relies purely on siloed data $(y)^{n}$.
KL Loss: We utilize the formulation given in Equation (8) to compute the KL loss as follows:
$L^{KL}=-\frac{1}{2}\Big{[}1+2\log(\sigma)-\mu^{2}-\sigma\Big{]}$ (8)
In Equation (8), the KL loss depends only on the latent space embeddings that result from the reparametrization trick described in Section III-A. Consequently, the backpropagation of the error resulting from the KL loss can be initiated from the server level itself. Therefore, the decomposed reconstruction and KL error terms can be leveraged to effect a computationally efficient training mechanism of the SplitVAEs framework.
@@ IV SplitVAEs Algorithmic Framework
There are two crucial characteristics of the SplitVAEs decomposition scheme that can be advantageous to its training. First, the KL loss, $L^{KL}$, is purely a factor of the latent space embeddings estimated by the server-based VAE framework. Second, the reconstruction error can be computed in a fully decentralized fashion by each of the edge locations locally without revealing the siloed datasets themselves. Thus, only the local reconstruction loss needs to be back-propagated from the edge-based AE decoder to the server-based VAE decoder. This step can be immediately followed by a backpropagation of the total loss, $L^{recon}+L^{KL}$, from the server-level VAE encoder to the edge-based AE encoders. As a result of our two-pronged backpropagation mechanism, we can enable the flow of insights back and forth between edge and server-level devices, leading to seamless learning of the overall spatiotemporal interdependencies among stakeholders.
A single training epoch in our proposed framework can be divided based on edge- and server-level algorithmic frameworks. In the following sections, we discuss in detail the decomposition of the respective frameworks about the forward and backward passes of the respective model components, i.e., the server-level VAE and the encoders and decoders of AE. Specifically, we consider a set of $\mathcal{N}$ edge nodes wherein $|\mathcal{N}|=N$. To accomplish the SplitVAEs training, we leverage several common distributed memory aggregation primitives such as Scatter and Gather, denoted as DM.Scatter and DM.Gather, respectively. Tensor.Concat, Tensor.Split, BACKPROP functions indicate tensor concatenation, split and backpropagation primitives respectively. We designate the server as the root process $r$, while the edge-level processes are denoted by $n$ similar to the concept of ranks in distributed memory frameworks. We use $k$ to denote iterative updates of the model parameters represented by $\theta^{n}_{E},\theta^{n}_{D},\phi,\theta$.
@@ IV-A Edge-based Algorithmic Framework
AE Encoder Forward Pass: The computational steps associated with the forward pass of the AE encoder are represented by the function EdgeEncFP. In EdgeEncFP, we consider local siloed training data $y^{b}_{e}$ which is passed through the AE encoder $F_{(\theta^{n}_{E})^{k}}$ to yield the low dimensional embedding $(x^{b}_{e})^{n}$. We utilize a DM.Gather operation to aggregate the embeddings at the server level.
@@
function EncEdgeFP(root:r,agent:n,epoch:$e$,batch:$b$)
@@
  DM.Gather(root:r,data:$(x^{b}_{e})^{n}$)
@@
function EdgeDecFP(root:r,agent:n,epoch:$e$,batch:$b$)
@@
AE Decoder Backward Pass: As an immediate next step, our implementation leverages EdgeDecBP to back-propagate the errors based on the reconstruction losses. In EdgeDecBP $(g_{\theta_{D}})^{b}_{e}=\nabla_{\theta_{D}}(L^{recon}_{n})^{b}_{e}$ represents the gradient of $(L^{recon}_{n})^{b}_{e}$ concerning $\theta_{D}$ and $\Delta(\tilde{x})^{b}_{e}$ represents the error at the input layer of the edge-based decoder. Finally, we use a DM.Gather operation to collect the errors of the low-dimensional embeddings $(\tilde{x}^{b}_{e})^{n}$ at the server.
@@
function EdgeDecBP(root:r,agent:n,epoch:$e$,batch:$b$)
  obtain $(\Delta\tilde{x})^{b}_{e},(g_{D})^{b}_{e}\leftarrow$ BACKPROP($G_{(\theta^{n}_{D})^{k}},(L^{BC}_{n})^{b}_{e}$)
@@
function EdgeEncBP(root:r,agent:n,epoch:$e$,batch:$b$)
@@
Algorithm 5 VAE Forward Pass
function VAEServerFP(root:r,epoch:$e$,batch:$b$)
  $\Big{[}(x^{b}_{e})^{1},(x^{b}_{e})^{2}\ldots(x^{b}_{e})^{N}\Big{]}\leftarrow$DM.Gather(root:r,data:NULL)
  acquire $x^{b}_{e}\leftarrow$ Tensor.Concat$\big{[}(x^{b}_{e})^{1},(x^{b}_{e})^{2}\ldots(x^{b}_{e})^{N}\big{]}$
  obtain $(\mu^{b}_{e},\sigma^{b}_{e})\leftarrow q_{\phi}(x^{b}_{e})$
  create $z^{b}_{e}\leftarrow\mu^{b}_{e}+\sigma^{b}_{e}\odot\epsilon^{b}_{e}$, where $\epsilon^{b}_{e}\sim\mathcal{N}(0,I)$
  estimate $\tilde{x}^{b}_{e}\leftarrow p_{\theta}(z^{b}_{e})$
  DM.Scatter(root:r,data:$\Big{[}(\tilde{x}^{b}_{e})^{1},(\tilde{x}^{b}_{e})^{2}\ldots(\tilde{x}^{b}_{e})^{N}\Big{]}$)
  compute KL loss $(L^{KL})^{b}_{e}$ using (8)
end function
@@ IV-B Server-based Algorithmic Framework
VAE Backward Pass: The backward pass of the VAE is represented by VAEServerBP. The backpropagated set of error values $(\Delta\tilde{x}^{b}_{e})^{1}\ldots(\Delta\tilde{x}^{b}_{e})^{N}$ resulting from $L^{recon}_{n}$ are acquired from each edge agent through DM.Gather. These error terms are concatenated and backpropagated through the probabilistic decoders to yield $(\Delta z)^{b}_{e},(g_{\theta})^{b}_{e}$ representing latent space error terms and probabilistic decoder gradients respectively. With a subsequent backpropagation of $(\Delta z)^{b}_{e}$ through the probabilistic encoder $q_{\phi}$, we obtain $(\Delta x^{BC})^{b}_{e},(g^{BC}_{\phi})^{b}_{e}$ which corresponds to the error terms at the input leaves of the VAE and the gradient values for the probabilistic encoder. It is important to note the errors and gradients accumulated so far on the VAE correspond only to the reconstruction error terms observed at the edge level. However, to complete VAE training we would also need to incorporate the KL loss. Since the KL loss is purely a function of the latent space embeddings $z^{b}_{e}$, we backpropagate $(L^{KL})^{b}_{e}$ computed during VAEServerFP and accumulate additional error values at the input leaves. The final set of accumulated errors $(\Delta x)^{b}_{e}$ are split and sent back to the corresponding edge agents using DM.Scatter.
@@ IV-C Decentralized Training Algorithm
Combining the training steps of the edge and server-level models provides us with Algorithm 7 and 8, for VAE and AE, respectively. In these algorithms, we abstract the training mechanism using the set of functions detailed in Sections IV-A and IV-B for a specific set of batch and epoch sizes.
@@
function VAEServerBP(root:r,epoch:$e$,batch:$b$)
  $\Big{[}(\Delta\tilde{x}^{b}_{e})^{1}\ldots(\Delta\tilde{x}^{b}_{e})^{N}\Big{]}\leftarrow$DM.Gather(root:r,data:NULL)
  $\Delta{\tilde{x}}^{b}_{e}\leftarrow$ Tensor.Concat$\big{[}(\Delta\tilde{x}^{b}_{e})^{1}\ldots(\Delta\tilde{x}^{b}_{e})^{N}\big{]}$
  $\triangleright$ Backpropagate BC loss
@@
  $[(\Delta\mu)^{b}_{e},(\Delta\sigma)^{b}_{e}]^{BC}\leftarrow$BACKPROP($R_{[\mu,\sigma]},(\Delta z)^{b}_{e}$)
  $(\Delta x^{BC})^{b}_{e},(g^{BC}_{\phi})^{b}_{e}\leftarrow$BACKPROP($q_{\phi},[(\Delta\mu)^{b}_{e},(\Delta\sigma)^{b}_{e}]^{BC}$)
  $\triangleright$ Backpropagate KL loss
  $[(\Delta\mu)^{b}_{e},(\Delta\sigma)^{b}_{e}]^{KL}\leftarrow$BACKPROP($R_{[\mu,\sigma]},{L_{KL}}^{b}_{e}$)
  $(\Delta x^{KL})^{b}_{e},(g^{KL}_{\phi})^{b}_{e}\leftarrow$BACKPROP($q_{\phi},[(\Delta\mu)^{b}_{e},(\Delta\sigma)^{b}_{e}]^{KL}$)
@@
  update $\theta^{k+1}\leftarrow\theta^{k}-\eta_{\theta}.(g_{\theta})^{b}_{e}$
  update $\phi^{k+1}\leftarrow\phi^{k}-\eta_{\phi}.\nabla_{\phi}((g^{BC}_{\phi})^{b}_{e}+(g^{KL}_{\phi})^{b}_{e})$
@@
  $(\Delta x)^{b}_{e}\leftarrow(\Delta x^{KL})^{b}_{e}+(\Delta x^{BC})^{b}_{e}$
@@
  $\Big{[}(\Delta{x}^{b}_{e})^{1}\ldots(\Delta{x}^{b}_{e})^{N}\Big{]}\leftarrow$Tensor.Split$\big{[}\Delta{x}^{b}_{e}\big{]}$
  DM.Scatter(root:r,data:$\Big{[}(\Delta{x}^{b}_{e})^{1},(\Delta{x}^{b}_{e})^{2}\ldots(\Delta{x}^{b}_{e})^{N}\Big{]}$
@@
Algorithm 7 Training Mechanism of VAE at Server-level
@@ V Benchmark Methods
In this section, we outline statistical benchmarking methods and metrics for our proposed SplitVAE framework.
@@ V-A Evaluation Methods
We compare the SplitVAEs with centralized scenario generation methods, such as the Gaussian copula [24, 25] and the Centralized-VAE method. Moreover, given that the datasets are multidimensional in nature, we untilize t-SNE plots [26] to visualize kernel density distributions of the generated scenarios, transformed into one-dimensional embeddings.
@@ V-B Evaluation Metrics
In addition to visual analyses, we also employ multiple quantitative evaluation metrics to evaluate the quality of the generated scenarios. Let $X\in\mathbb{R}^{m_{1}\times d}$ be the observed data and $Y\in\mathbb{R}^{m_{2}\times d}$ be the generated scenarios, where $m_{i}\in\{1,2\}$ is the number of observations and $d$ is the number of features. Lower values indicate better model performance for each of the metrics discussed below. Thus, $X$ and $Y$ can be expressed as: $X=x_{1},x_{2},...,x_{m_{1}}$, where $x_{i}\in\mathbb{R}^{d}$ for all $i=\{1,2,...,m_{1}\}$; $Y=y_{1},y_{2},...,y_{m_{2}}$, where $y_{i}\in\mathbb{R}^{d}$ for all $i=\{1,2,...,m_{2}\}$. The metrics are as follows:
Fréchet inception distance (FID): The FID score is usually employed to evaluate deep generative models (DGMs) [27, 28]. The FID measures the Wasserstein distance between two probability distributions $f_{X}(x)$ and $f_{Y}(y)$. Since the observed data, $X$, and generated scenarios, $Y$, are multivariate Gaussian distributions, i.e, $X\sim\mathcal{N}(\mu_{x},\Sigma_{x})$ and $Y\sim\mathcal{N}(\mu_{y},\Sigma_{y})$, the FID is given as [29]:
@@
Continuous ranked probability score: The CRPS [31] measures the squared distance between a single multivariate observed observation $y$ and the median of $F(x)$ which is the empirical cumulative distribution of generated scenarios [32]. For an $m_{2}$ number of observations, the CRPS is given as:
@@
Thus, the evaluation methods and metrics form an important component of our experiment strategy enabling us to analyze the quality of generated scenarios from multiple facets.

### v2

@@ III-D Loss Functions
Reconstruction Loss: We employ the binary cross-entropy (BC) loss function at the $n$-th edge location for all $n\in\mathcal{N}$ to measure the mean reconstruction error between predicted values $\tilde{y}^{n}$ and observed values $y^{n}$ as represented in Equation (7).
$L^{\text{BC}}_{n}=-\frac{1}{B}\sum\limits_{b=1}^{B}\Big{[}\tilde{y}^{b}log(y^{b})+(1-\tilde{y}^{b})log(1-y^{b})\Big{]}$ (7)
In Equation (7), let $B,\tilde{y}_{b},y_{b}$ be the batch size, predicted, and target data for scenario generation, respectively. To preserve the simplicity of notations, we assume without loss of generality that the loss function $L^{\text{BC}}_{n}$ is computed with respect to flattened representations of local tensor data and estimates both $y^{n}$ and $\tilde{y}^{n}$ of dimension $d^{n}$. We note that the computation of $L^{\text{BC}}_{n}$ can be carried out independently at each edge location since it relies purely on siloed data, represented by $y^{n}$.
KL Loss: We utilize the formulation given in Equation (8) to derive the relevant KL loss function as follows:
$L^{\text{KL}}=-\frac{1}{2}\Big{[}1+2\log(\hat{\sigma})-\hat{\mu}^{2}-\hat{\sigma}\Big{]}$ (8)
In Equation (8), the computed KL loss depends only on the latent space embeddings resulting from the reparametrization trick described in Section III-A. Consequently, the backpropagation of the error resulting from the KL loss can be initiated from the server level itself. Therefore, the decomposed reconstruction and KL error terms can be leveraged to effect a computationally efficient training mechanism of the SplitVAEs framework.
@@ IV SplitVAEs Algorithmic Framework
There are two crucial characteristics of the SplitVAE decomposition scheme that can be advantageous to its training. First, the KL loss, $L^{\text{KL}}$, is purely a factor of the latent space embeddings estimated by the server-based VAE framework. Second, the reconstruction error can be computed in a fully decentralized fashion by each of the edge locations locally without revealing the siloed datasets themselves. Thus, only the local reconstruction loss needs to be back-propagated from the edge-based AE decoder to the server-based VAE decoder. This step can be immediately followed by a backpropagation of the total loss, $L^{\text{recon}}+L^{\text{KL}}$, from the server-level VAE encoder to the edge-based AE encoders. As a result of our two-pronged backpropagation mechanism, we can enable the flow of insights back and forth between edge and server-level devices, leading to seamless learning of the overall spatiotemporal interdependencies among stakeholders.
A single training epoch in our proposed framework can be divided based on edge and server level algorithmic frameworks. In the following sections, we discuss the decomposition of the forward and backward passes of the server-level VAE as well as the encoders and decoders of edge-level AEs. Specifically, without loss of generality, we consider a set of $\mathcal{N}$ edge nodes wherein $|\mathcal{N}|=N$. To accomplish the training of SplitVAEs, we leverage several common distributed memory aggregation primitives such as the scattering and gathering operations, denoted by DM.Scatter and DM.Gather, respectively. Let Tensor.Concat, Tensor.Split, and BACKPROP functions indicate tensor concatenation, split and backpropagation primitives, respectively. We designate the server as the root process $r$, while the edge-level processes are denoted by $n$ and are similar to the concept of ranks in distributed memory frameworks. We use $k$ to denote iterative updates of the model parameters represented by $\theta^{n}_{E},\theta^{n}_{D},\phi,\theta$.
@@ IV-A Edge-based Algorithmic Framework
AE Encoder Forward Pass: The computational steps associated with the forward pass of the AE encoder are represented by the function EdgeEncFP. In EdgeEncFP, we consider local siloed training data $y^{b}_{e}$ that is passed through the AE encoder $F_{(\theta^{n}_{E})^{k}}$ to yield the low dimensional embedding $(x^{b}_{e})^{n}$. We utilize a DM.Gather operation to aggregate the embeddings at the server level.
@@
function EncEdgeFP(root:r, agent:n, $e$, $b$)
@@
  Execute DM.Gather(root:r, data:$(x^{b}_{e})^{n}$)
@@
function EdgeDecFP(root:r, agent:n, $e$, $b$)
@@
AE Decoder Backward Pass: As an immediate next step, our implementation leverages EdgeDecBP to back-propagate the errors based on the reconstruction losses. In EdgeDecBP $(g_{\theta_{D}})^{b}_{e}=\nabla_{\theta_{D}}(L^{\text{recon}}_{n})^{b}_{e}$ represents the gradient of $(L^{\text{recon}}_{n})^{b}_{e}$ concerning $\theta_{D}$ and $\Delta(\tilde{x})^{b}_{e}$ represents the error at the input layer of the edge-based decoder. Finally, we use a DM.Gather operation to collect the errors of the low-dimensional embeddings $(\tilde{x}^{b}_{e})^{n}$ at the server.
@@
function EdgeDecBP(root:r, agent:n, $e$, $b$)
  Obtain $(\Delta\tilde{x})^{b}_{e}$
  Obtain $(g_{D})^{b}_{e}\leftarrow$ BACKPROP($G_{(\theta^{n}_{D})^{k}},(L^{{}^{\text{BC}}}_{n})^{b}_{e}$)
@@
function EdgeEncBP(root:r, agent:n, $e$, $b$)
@@ IV-B Server-based Algorithmic Framework
VAE Backward Pass: The backward pass of the VAE is represented by VAEServerBP. The backpropagated set of error values, denoted by $\Big{[}(\Delta\tilde{x}^{b}_{e})^{1},(\Delta\tilde{x}^{b}_{e})^{2},\ldots,(\Delta\tilde{x}^{b}_{e})^{N}\Big{]}$ resulting from $L^{\text{recon}}_{n}$ are acquired from each edge agent through DM.Gather. These error terms are concatenated and back-propagated through the probabilistic decoders to yield $(\Delta z)^{b}_{e}$ and $(g_{\theta})^{b}_{e}$ representing latent space error terms and probabilistic decoder gradients, respectively. With a subsequent backpropagation of $(\Delta z)^{b}_{e}$ through the probabilistic encoder $q_{\phi}$, we obtain $(\Delta x^{{}^{\text{BC}}})^{b}_{e}$ and $(g^{{}^{\text{BC}}}_{\phi})^{b}_{e}$ corresponding to the error terms at the input leaves of the VAE and the gradient values for the probabilistic encoder, respectively. More importantly, the errors and gradients accumulated so far on the VAE correspond to the reconstruction error terms observed at the edge level only.
Algorithm 5 VAE Forward Pass
function VAEServerFP(root:r, $e$, $b$)
  $\Big{[}(x^{b}_{e})^{1},\ldots,(x^{b}_{e})^{N}\Big{]}\leftarrow$ DM.Gather(root:r,data:NULL)
  Acquire $x^{b}_{e}\leftarrow$ Tensor.Concat$\big{[}(x^{b}_{e})^{1},(x^{b}_{e})^{2},\ldots,(x^{b}_{e})^{N}\big{]}$
  Obtain $(\hat{\mu}^{b}_{e},\hat{\sigma}^{b}_{e})\leftarrow q_{\phi}(x^{b}_{e})$
  Create $z^{b}_{e}\leftarrow\hat{\mu}^{b}_{e}+\hat{\sigma}^{b}_{e}\odot\epsilon^{b}_{e}$, where $\epsilon^{b}_{e}\sim\mathcal{N}(0,I_{s})$
  Estimate $\tilde{x}^{b}_{e}\leftarrow p_{\theta}(z^{b}_{e})$
  DM.Scatter(root:r, data:$\Big{[}(\tilde{x}^{b}_{e})^{1},\ldots,(\tilde{x}^{b}_{e})^{N}\Big{]}$)
  Compute KL loss $(L^{\text{KL}})^{b}_{e}$ using (8)
end function
However, to complete VAE training, we would also need to incorporate the KL loss. Since the KL loss is purely a function of the latent space embeddings $z^{b}_{e}$, we backpropagate $(L^{\text{KL}})^{b}_{e}$ computed during VAEServerFP and accumulate additional error values at the input leaves. The final set of accumulated errors $(\Delta x)^{b}_{e}$ are split and sent back to the corresponding edge agents using DM.Scatter.
@@ IV-C Decentralized Training Algorithm
Combining the training steps of the edge and server-level models provides us with Algorithm 7 and 8 for VAE and AE, respectively. In these algorithms, we abstract the training mechanism using the set of functions detailed in Sections IV-A and IV-B for a specific set of batch and epoch sizes.
@@
function VAEServerBP(root:r, $e$, $b$)
  $\Big{[}(\Delta\tilde{x}^{b}_{e})^{1},\ldots,(\tilde{x}^{b}_{e})^{N}\Big{]}\leftarrow$DM.Gather(root:r, data:NULL)
  $\Delta{\tilde{x}}^{b}_{e}\leftarrow$ Tensor.Concat$\big{[}(\Delta\tilde{x}^{b}_{e})^{1},\ldots,(\Delta\tilde{x}^{b}_{e})^{N}\big{]}$
  $\triangleright$ Backpropagate BC loss:
@@
  $\Big{[}(\Delta\hat{\mu})^{b}_{e},(\Delta\hat{\sigma})^{b}_{e}\Big{]}^{\text{BC}}\leftarrow$BACKPROP($R_{\hat{\mu},\hat{\sigma}},(\Delta z)^{b}_{e}$)
  $(\Delta x^{{}^{\text{BC}}})^{b}_{e},(g^{{}^{\text{BC}}}_{\phi})^{b}_{e}\leftarrow$BACKPROP($q_{\phi},\Big{[}(\Delta\hat{\mu})^{b}_{e},(\Delta\hat{\sigma})^{b}_{e}\Big{]}^{\text{BC}}$)
  $\triangleright$ Backpropagate KL loss:
  $[(\Delta\hat{\mu})^{b}_{e},(\Delta\hat{\sigma})^{b}_{e}]^{\text{KL}}\leftarrow$ BACKPROP($R_{\hat{\mu},\hat{\sigma}},{L_{\text{KL}}}^{b}_{e}$)
  $(\Delta x^{\text{KL}})^{b}_{e},(g^{\text{KL}}_{\phi})^{b}_{e}\leftarrow$BACKPROP($q_{\phi},\Big{[}(\Delta\hat{\mu})^{b}_{e},(\Delta\hat{\sigma})^{b}_{e}\Big{]}^{\text{KL}}$)
@@
  $\theta^{k+1}\leftarrow\theta^{k}-\eta_{\theta}.(g_{\theta})^{b}_{e}$
  $\phi^{k+1}\leftarrow\phi^{k}-\eta_{\phi}.\nabla_{\phi}((g^{{}^{\text{BC}}}_{\phi})^{b}_{e}+(g^{\text{KL}}_{\phi})^{b}_{e})$
@@
  $(\Delta x)^{b}_{e}\leftarrow(\Delta x^{\text{KL}})^{b}_{e}+(\Delta x^{{}^{\text{BC}}})^{b}_{e}$
@@
  $\Big{[}(\Delta{x}^{b}_{e})^{1},\ldots,(\Delta{x}^{b}_{e})^{N}\Big{]}\leftarrow$ Tensor.Split$\big{[}\Delta{x}^{b}_{e}\big{]}$
  DM.Scatter(root:r, data:$\Big{[}(\Delta{x}^{b}_{e})^{1},(\Delta{x}^{b}_{e})^{2},\ldots,(\Delta{x}^{b}_{e})^{N}\Big{]}$
@@
Algorithm 7 Training Mechanism of Server-based VAE
@@ V Benchmark Methods
In this section, we discuss the relevant benchmarking strategies for SplitVAEs by outlining the evaluation methods as well as the evaluation metrics employed in this study.
@@ V-A Evaluation Methods
We compare SplitVAEs with centralized scenario generation methods, including the Gaussian copula [25, 26] and the Centralized-VAE. The Gaussian copula models the joint probability distribution function (PDF) to capture the interdependent structures, specifically the spatiotemporal interdependencies within the data. Meanwhile, in SplitVAEs and Centralized-VAE, these interdependencies are modeled within the latent space $z$ when training the probabilistic encoder $q_{\phi}(z|x)$ and decoder $p_{\theta}(x|z)$ pair, as mentioned in Section III. Further, given that the datasets and their corresponding scenarios are multidimensional in nature, we employ t-SNE graphs[27] to visualize kernel density distributions of the observed and generated data, which are transformed into one-dimensional embeddings.
@@ V-B Evaluation Metrics
In addition to visual analyses, we employ multiple quantitative evaluation metrics to evaluate the quality of the generated scenarios. For each metric, lower values indicate a better model performance.
Let $X\in\mathbb{R}^{m_{1}\times d}$ be the real data and $Y\in\mathbb{R}^{m_{2}\times d}$ be the generated scenarios, each comprising of $d$ features with $m_{1},m_{2}$ number of observations respectively. Thus, $X$ and $Y$ can be expressed as: $X=\{x_{1},x_{2},...,x_{m_{1}}\}$, where $x_{i}\in\mathbb{R}^{d}$ for all $1\leq i\leq m_{1}$; $Y=\{y_{1},y_{2},...,y_{m_{2}}\}$, where $y_{i}\in\mathbb{R}^{d}$ for all $1\leq i\leq m_{2}$. Hence, we can derive each metric as follows:
Fréchet inception distance (FID): The FID score is a widely accepted, state-of-the-art metric employed to evaluate deep generative models (DGMs) [28, 29]. The FID measures the Wasserstein distance between two PDFs $f_{X}(x)$ and $f_{Y}(y)$. In our study, the observed dataset and the generated scenarios, have finite values pertaining to their respective mean and covariance matrices denoted by $\mu_{x},\mu_{y}$ and $\Sigma_{x},\Sigma_{y}$ respectively. Under such conditions, based on [30] the FID can be defined as follows:
@@
Continuous ranked probability score: The CRPS [32] measures the squared distance between a single multivariate observed observation $y$ and the median of $F(x)$ which is the empirical cumulative distribution of generated scenarios [33]. For $m_{2}$ number of observations, the CRPS is given as: